# Cost-aware Photonic Graph State Generation: A Graphical Framework

### v1

@@ VII.3.5 Extra emitters
Furthermore, as demonstrated in the case analysis, the number of costly operations in each step depend on the size and overlap of the sets $\mathcal{J}$, $\mathcal{K}$, and $\mathcal{M}$. As seen in eqs. 22, 24 and 26, the composition of these sets is fully determined in terms of emitter rows. Since there is a corresponding emitter row for each active emitter in the system, the use of extra emitters allows one to have flexibility in making up the sets $\mathcal{J}$, $\mathcal{K}$, and $\mathcal{M}$, to minimize the number of costly operations in each step.
@@ VIII.1 Necessity of Condition I
$\exists\,\tau\in\mathcal{T}^{*}\;\text{such that}\;G^{\prime}(n)\xrightarrow{\tau}G^{\prime}(N)=G,$ (28)
@@
$\text{rank}\,[B^{\prime}(n)]<\text{rank}\,[B(n)],$ (29)
@@
$\text{rank}\,[B^{\prime}(n)]\geq\text{rank}\,[B_{N}^{\prime}(n)],$ (30)
@@
$\text{rank}\,[B^{\prime}(n)]\geq\text{rank}\,[B(n)]$
@@
$-\text{Tr}(\rho_{\mathcal{A}}\log\rho_{\mathcal{A}})\,$, (32)
@@
$\mathcal{I}_{\mathcal{A}}=|\mathcal{A}|+\text{rank}(B).$ (37)
@@
$S(\rho_{\mathcal{A}})=\mathcal{I}_{\mathcal{A}}-|\mathcal{A}|=\text{rank}(B).$ (38)
@@
Any allowed sequence of operations on a graph state during its generation can be viewed as a series of LC operations divided by two‑qubit interactions between emitters, photon emission events, or measurements. Since we allow no two-qubit gates on emitted photons, the edge creation/removal between the two subsystems of the state is only possible via applying local Clifford operations, whose effects on the edge structure of a graph state are shown to be reproducible with a sequence of local complementation (LC) transformations. Therefore, by using lemma 1, one can deduce that such local operations cannot change the rank of a biadjacency matrix. The allowed two-qubit gates do not change the rank either since they are limited to only the second subsystem and create no edge across the bipartition, thus leaving the biadjacency matrix unchanged. Besides, photons emitted in future would not be a member of the first subsystem and are each introduced to the second subsystem as a leaf qubit attached to one emitter (see FIG. 1 c), thus not affecting the rank of the biadjacency matrix. Lastly, it is evident that measurements cannot increase the bipartite entanglement either. Having proved that rank cannot increase under the allowed operations for a fixed bipartition, one can use theorem 1, relating rank to entanglement entropy, and complete the proof of the proposition. ∎
@@ VIII.2 Necessity of Condition II
$\begin{split}\mathcal{N}^{\mathrm{ex}}(L)&=\mathcal{N}(L)-\{P\},\\ \mathcal{N}^{\mathrm{ex}}(P)&=\mathcal{N}(P)-\{L\},\end{split}$ (40)
@@
where $\{C_{j}(n)\}$ and $\{C^{\prime}_{j}(n)\}$ are the set of columns of $B(n)$ and $B^{\prime}(n)$ respectively. We are now in a place to demonstrate that the specific connectivity configuration between the emitters and photon required by condition (II) is necessary for the statement 41 to hold.
@@
$\alpha^{l}_{k}=1\Leftrightarrow[B^{T}_{\mathrm{echelon}}]_{k\,l}=1,$ (46)
@@
$\beta^{j}_{k}=\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}.$ (53)
@@
$\beta^{i}_{k}=\sum_{j\in S(i)}\beta^{j}_{k}.$ (54)
By combining eqs. 53 and 54 we get:
$\beta^{i}_{k}=\sum_{j\in S(i)}\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}.$ (55)
@@
$\beta^{i}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k}.$ (56)
@@
$\sum_{j\in S(i)}\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k},$ (57)
@@
$=\sum_{i^{\prime}<m}\left(\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\right)\beta^{i^{\prime}}_{k}.$ (59)
@@
$\sum_{i^{\prime}<m}\left(\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\right)\beta^{i^{\prime}}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k}=\beta^{i}_{k}.$ (60)
@@
$\alpha^{i}_{i^{\prime}}=\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\quad\text{for all }i^{\prime}<m.$ (61)
@@
$\alpha^{i}_{k}=\sum_{j\in S(i)}\alpha^{j}_{k}\quad\text{for all }k<m.$ (62)
@@
$R_{i}=\sum_{j<m}\alpha^{i}_{j}R_{e_{j}}.$ (63)
Following the same reasoning used between eqs. 45 and 51, the elements of $B^{\prime}$ can be written as:
@@
where we used eqs. 46, 47 and 48 in order to get to the last biconditional. Here the set $S(i)$ is defined by the expansion:
$R_{i}=\sum_{j\in S(i)}R_{j},$ (65)
@@ VIII.3 Necessity of Condition III
$G^{\prime}(n)\xrightarrow{\tau}G,\quad\tau\in\mathcal{T}^{*}$
@@
$G^{\prime}(n)\xrightarrow{\tau_{1}}H\xrightarrow{\tau_{2}}G,$ (69)
@@
$\forall\,\tilde{H},\ \exists\,\tilde{G}\in\mathrm{PLC}(G,\mathcal{P})\;\text{such that}\ \tilde{G}[\mathcal{P}]=\tilde{H}[\mathcal{P}].$ (70)
@@
$H\xrightarrow{\text{LC on $h$}}\tilde{H}$,
@@
$\tilde{G}[\mathcal{P}]=\tilde{H}[\mathcal{P}].$
@@
$\exists\,\tilde{G}\in\mathrm{PLC}(G,\mathcal{P})\;\text{such that}\ \tilde{G}[\mathcal{P}]={G^{\prime}(n)}[\mathcal{P}].$
@@ Appendix C Edge-density Dependence in Random Graphs
Because the cnot count generally increases with edge density, we utilize the method introduced in Ref. [49], which replaces the target graph with a locally equivalent alternative having fewer edges, thereby reducing the cnot cost. To identify such alternatives, we use the deterministic edge reduction strategy described in Appendix C of Ref. [49], which is an efficient process that finishes in at most $\mathcal{O}(N^{2})$ steps. This edge reduction is applied as a built-in first step before executing the main algorithm to construct a generation circuit. Figure 24 (b) shows the resulting cnot suppression, shifting the peak from approximately 90% edge density to around 50%, and improving performance over a broader range of $p$ values.

### v2

@@ VII.3.5 Extra emitters
Furthermore, as demonstrated in the case analysis, the number of costly operations in each step depend on the size and overlap of the sets $\mathcal{J}$, $\mathcal{K}$, and $\mathcal{M}$. As seen in eqs. 23, 25 and 27, the composition of these sets is fully determined in terms of emitter rows. Since there is a corresponding emitter row for each active emitter in the system, the use of extra emitters allows one to have flexibility in making up the sets $\mathcal{J}$, $\mathcal{K}$, and $\mathcal{M}$, to minimize the number of costly operations in each step.
VII.3.6 Emission order
Scheduling the emission photons is a common practice to avoid unnecessary delay lines and photon storage. If the intended application of the generated state does not require a fixed emission order, the emission sequence can be treated as an additional degree of freedom with potentially significant effects on the generation circuit. Changing the ordering of nodes in a graph alters the corresponding adjacency matrix and the linear dependency relations between its rows and columns. As a consequence, the eligibility conditions (I) and (II), which depend on these relations, are generally not invariant under node reordering. One immediate implication is a possible change in the number of emitters required for state generation. This can be exploited to identify generation circuits requiring fewer emitters and, potentially, a reduced number of emitter–emitter two-qubit gates when a search over different emission orders is performed. However, since the search space grows factorially with graph size, such an approach does not constitute a scalable optimization strategy. Indeed, Ref. [38] argues that, in general, finding the optimal emission order that minimizes the number of emitters is an NP-hard problem. Nevertheless, heuristic methods can be employed to select suitable emission orders in specific cases. For example, for loop-free graphs (tree graphs), ordering nodes according to a depth-first traversal maximizes the continuity of entanglement along each branch, often resulting in a reduced emitter requirement. This strategy is the default ordering used to obtain the tree-graph results presented in Table 1. Moreover, for parameterized structured graphs, one may perform an exhaustive search over small instances to identify optimal orderings and subsequently generalize the observed patterns to larger cases. A similar approach was adopted in Ref. [23].
@@ VIII.1 Necessity of Condition I
$\exists\,\tau\in\mathcal{T}^{*}\;\text{such that}\;G^{\prime}(n)\xrightarrow{\tau}G^{\prime}(N)=G$ (29)
@@
$\text{rank}\,[B^{\prime}(n)]<\text{rank}\,[B(n)]$ (30)
@@
$\text{rank}\,[B^{\prime}(n)]\geq\text{rank}\,[B_{N}^{\prime}(n)]$ (31)
@@
$\text{rank}\,[B^{\prime}(n)]\geq\text{rank}\,[B(n)]$ (32)
@@
$-\text{Tr}(\rho_{\mathcal{A}}\log\rho_{\mathcal{A}})\,$ (34)
@@
$\mathcal{I}_{\mathcal{A}}=|\mathcal{A}|+\text{rank}(B)$ (39)
@@
$S(\rho_{\mathcal{A}})=\mathcal{I}_{\mathcal{A}}-|\mathcal{A}|=\text{rank}(B)$ (40)
@@
Any allowed sequence of operations on a graph state during its generation can be viewed as a series of LC operations divided by two‑qubit interactions between emitters, photon emission events, or measurements. Since we allow no two-qubit gates on emitted photons, the edge creation/removal between the two subsystems of the state is only possible via applying local Clifford operations, whose effects on the edge structure of a graph state are shown to be reproducible with a sequence of local complementation (LC) transformations. Therefore, by using lemma 1, one can deduce that such local operations cannot change the rank of a biadjacency matrix. The allowed two-qubit gates do not change the rank either since they are limited to only the second subsystem and create no edge across the bipartition, thus leaving the biadjacency matrix unchanged. Besides, photons emitted in future would not be a member of the first subsystem and are each introduced to the second subsystem as a leaf qubit attached to one emitter, as seen in Fig. 1(c), thus not affecting the rank of the biadjacency matrix. Lastly, it is evident that measurements cannot increase the bipartite entanglement either. Having proved that rank cannot increase under the allowed operations for a fixed bipartition, one can use theorem 1, relating rank to entanglement entropy, and complete the proof of the proposition. ∎
@@ VIII.2 Necessity of Condition II
$\begin{split}\mathcal{N}^{\mathrm{ex}}(L)&=\mathcal{N}(L)-\{P\}\\ \mathcal{N}^{\mathrm{ex}}(P)&=\mathcal{N}(P)-\{L\}\end{split}$ (42)
@@
where $\{C_{j}(n)\}$ and $\{C^{\prime}_{j}(n)\}$ are the set of columns of $B(n)$ and $B^{\prime}(n)$ respectively. We are now in a place to demonstrate that the specific connectivity configuration between the emitters and photon required by condition (II) is necessary for the statement 43 to hold.
@@
$\alpha^{l}_{k}=1\Leftrightarrow[B^{T}_{\mathrm{echelon}}]_{k\,l}=1$ (48)
@@
$\beta^{j}_{k}=\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}$ (55)
@@
$\beta^{i}_{k}=\sum_{j\in S(i)}\beta^{j}_{k}$ (56)
By combining eqs. 55 and 56 we get:
$\beta^{i}_{k}=\sum_{j\in S(i)}\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}$ (57)
@@
$\beta^{i}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k}$ (58)
@@
$\sum_{j\in S(i)}\sum_{i^{\prime}<m}\alpha^{j}_{i^{\prime}}\beta^{i^{\prime}}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k}$ (59)
@@
$=\sum_{i^{\prime}<m}\left(\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\right)\beta^{i^{\prime}}_{k}$ (61)
@@
$\sum_{i^{\prime}<m}\left(\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\right)\beta^{i^{\prime}}_{k}=\sum_{i^{\prime}<m}\alpha^{i}_{i^{\prime}}\beta^{i^{\prime}}_{k}=\beta^{i}_{k}$ (62)
@@
$\alpha^{i}_{i^{\prime}}=\sum_{j\in S(i)}\alpha^{j}_{i^{\prime}}\quad\text{for all }i^{\prime}<m$ (63)
@@
$\alpha^{i}_{k}=\sum_{j\in S(i)}\alpha^{j}_{k}\quad\text{for all }k<m$ (64)
@@
$R_{i}=\sum_{j<m}\alpha^{i}_{j}R_{e_{j}}$ (65)
Following the same reasoning used between eqs. 47 and 53, the elements of $B^{\prime}$ can be written as:
@@
where we used eqs. 48, 49 and 50 in order to get to the last biconditional. Here the set $S(i)$ is defined by the expansion:
$R_{i}=\sum_{j\in S(i)}R_{j}$ (67)
@@ VIII.3 Necessity of Condition III
$G^{\prime}(n)\xrightarrow{\tau}G,\quad\tau\in\mathcal{T}^{*}$ (71)
@@
$G^{\prime}(n)\xrightarrow{\tau_{1}}H\xrightarrow{\tau_{2}}G$ (72)
@@
$\forall\,\tilde{H},\ \exists\,\tilde{G}\in\mathrm{PLC}(G,\mathcal{P})\;\text{such that}\ \tilde{G}[\mathcal{P}]=\tilde{H}[\mathcal{P}]$ (73)
@@
$H\xrightarrow{\text{LC on $h$}}\tilde{H}$
@@
$\tilde{G}[\mathcal{P}]=\tilde{H}[\mathcal{P}]$ (75)
@@
$\exists\,\tilde{G}\in\mathrm{PLC}(G,\mathcal{P})\;\text{such that}\ \tilde{G}[\mathcal{P}]={G^{\prime}(n)}[\mathcal{P}]$ (76)
@@ Appendix C Edge-density Dependence in Random Graphs
Because the cnot count generally increases with edge density, we utilize the method introduced in Ref. [23], which replaces the target graph with a locally equivalent alternative having fewer edges, thereby reducing the cnot cost. To identify such alternatives, we use the deterministic edge reduction strategy described in Appendix C of Ref. [23], which is an efficient process that finishes in at most $\mathcal{O}(N^{2})$ steps. This edge reduction is applied as a built-in first step before executing the main algorithm to construct a generation circuit. Figure 25(b) shows the resulting cnot suppression, shifting the cnot cost peak from approximately 80% edge density to around 50%, and improving performance over a broader range of $p$ values.
@@ Appendix D Runtime
C: encoded 6-ring graph elementary circuit blocks
The parity-encoded 6-ring graphs are obtained by placing six logical nodes, each comprised of $m\times n$ physical qubits, in a ring configuration. The quantum circuit generating the state contains repeated circuit blocks due to the structural symmetries of these graphs and the repeated form of the entanglement within each logical node. First, let us employ the algorithm “time-reversed-2” to determine the generation circuit. In Fig. S2, we show the elementary circuit blocks that are repeated $n$ times to generate each logical qubit. Each block includes $m$ photon emissions, yielding a total of $n\times m$ qubits per logical node. From the number of two-qubit gates used in each block, one can determine the overall scaling of two-qubit gate usage with graph size. Explicitly, in this case, $1,2,1,2,1,$ and $1$ CNOTs are used in the elementary circuit blocks corresponding to the six logical qubits, respectively. Accounting for the $n$ repetitions, the total scaling is $8n$. Note that the generation circuit can have a different structure when transitioning between logical nodes; such transition blocks introduce an offset in the exact number of two-qubit gates used in the full circuit. We have analyzed the generation circuits for a range of values of $n$ and $m$, summarized in Table S1, and find that the exact number of CNOTs is 8n48n-4 for any $n,m>2$. This is consistent with the expected $8n$ scaling in the asymptotic large-$n$ regime. Note that an increase in $m$ only affects the number of times the emission operations are repeated in each block and does not result in more emitter-emitter CNOTs.
We next consider the new “Graph Builder” algorithm, whose corresponding elementary circuit blocks are shown in Fig. S3. In this construction, all six logical nodes are generated by circuit blocks that each contain a single two-qubit gate. Repeating these blocks $n$ times therefore leads to a total two-qubit gate scaling of $6n$. As before, transition blocks between logical nodes introduce a constant offset in the exact gate count. From the explicit generation circuits evaluated for a range of values of $n$ and $m$, summarized in Table S1, we find that the total number of CZ gates is given by $6n+4$. This again agrees with the expected $6n$ scaling in the asymptotic large-$n$ limit.
Finally, we consider the “time-reversed-1” algorithm. The generation circuit in this case can still be understood in terms of repeated circuit structures and the corresponding elementary blocks are presented in Fig. S4, showing the use of $2,2,1,2,1$, and $1$ CNOTs for each of the six logical node circuit blocks respectively. The overall scaling is thus $9n$. The exact two-qubit gate cost can also be extracted from the full generation circuits, with the results summarized in Table S1. We find that the total number of two-qubit gates scales as 9n79n-7, confirming our prediction based on the circuit blocks.
Figure S2: Elementary circuit blocks used in the generation of the parity-encoded 6-ring graph using “time-reversed-2” algorithm. Three distinct circuit blocks are shown, corresponding to the logical nodes $(1,6)$, $(2,4)$, and $(3,5)$, as indicated above each block. Each block is repeated $n$ times to generate the corresponding logical node. Photon emissions are to be repeated $m$ times in each block and are represented by arrows originated from the emitters in the quantum circuit. Note that the first block and the third block are related by a symmetry, namely relabeling of the three emitters. However, the second block looks different from other two blocks and involves two (instead of one) CNOT gates.
Table S1: Number of two-qubit gates used for the generation of on 6-ring parity encoded graphs for different $(n,m)$ encoding configurations. The algorithms TR1, and TR2, are referring to the two time-reversed generation algorithms defined in the main text. The number of emitters required ($n_{e}$), number of nodes (NN), and number of edges ($|E|$) in each graph are also reported.
Encoding parameters $(n,m)$ $n_{e}$ NN $\mid E\mid$ Graph Builder TR2 TR1
(4,2) 3 48 78 28 28 29
(4,3) 3 72 144 28 28 29
(5,3) 3 90 174 34 36 38
(6,3) 3 108 204 40 44 47
(5,4) 3 120 264 34 36 38
(6,4) 3 144 306 40 44 47
(7,4) 3 168 348 46 52 56
(10,4) 3 240 474 64 76 83
(9,6) 3 324 744 58 68 73
(17,4) 3 408 768 106 132 146
(12,7) 3 504 1152 76 92 101
Figure S3: Elementary circuit blocks used in the generation of the parity-encoded 6-ring graph using the “Graph Builder” algorithm. Four distinct circuit blocks are shown, corresponding to the logical nodes $(1)$, $(2,4)$, $(3,5)$, and $(6)$, as indicated above each block. Each block is repeated $n$ times to generate the corresponding logical node. Photon emissions are to be repeated $m$ times in each block (the shown example is for $m=2$) and are represented by arrows originated from the emitters in the quantum circuit. It worth noting that all four blocks are equivalent under a permutation of emitter qubit labels.
Figure S4: Elementary circuit blocks used in the generation of the parity-encoded 6-ring graph using the “time-reversed-1” algorithm. Three distinct circuit blocks are shown, corresponding to the logical nodes $(1,2,4)$, $(3,5)$, and $(6)$, as indicated below each block. Each block is repeated $n$ times to generate the corresponding logical node. Photon emissions are to be repeated $m$ times in each block (the shown example is for $m=2$) and are represented by arrows originated from the emitters in the quantum circuit. Note that the second and third blocks are the same up to a relabeling of emitters. In contrast, the first block involves an extra CNOT.
D: Explicit example for leveraging additional emitters
In this section, we present a simple and explicit example illustrating the potential advantage of allowing additional emitters in the construction of photonic graph-state generation circuits. We first apply the default Graph Builder algorithm that constrains the construction to the minimum number of emitters required for the target graph. We then compare this with an alternative approach in which this constraint is relaxed and additional emitters are permitted. The introduction of extra emitters follows a greedy strategy; before the emission of each photon, the option of introducing a new emitter is evaluated and is accepted only if it reduces the number of two-qubit gates required to prepare the system for the emission of the next immediate photon. Note that this strategy is local in nature and more sophisticated approaches, for example, those involving multi-step look-ahead, may yield further improvements. The target graph is a random 8 node graph shown in Fig. S5 and the resulting generation circuits are shown in Fig. S6. Due to the large circuit size, and to maintain visual clarity, we also provide a truncated version of each circuit showing all gates up to the final occurrence of a two-qubit gate; see Fig. S7. Allowing additional emitters leads to a reduction in the number of emitter–emitter CZ gates from 7 to 5, as well as a reduction in the two-qubit depth of the circuit from 7 to 4. Here, the two-qubit depth is defined in terms of the logical dependency structure of the circuit and is extracted from the directed acyclic graph (DAG) representation of the quantum circuit, considering emitters from their initializations to subsequent measurements.
Figure S5: The target graph state.
Figure S6: The generation circuits. (a) Using minimum number of emitters. (b) Allowing extra emitters. The advantage is clear in both number of two-qubit operations—highlighted in the circuit—and the depth of the circuit.
Figure S7: A truncated version of circuits in Fig. S6.